OpenStudy (anonymous):

for the given function f and g . find the requested composite function value. f(x)=sqrt (x+1) g(x)= 3x find (f o g)(1) please give me the answer..

OpenStudy (anonymous):

i have f((g(1))= square root (3+1) = square root (4) =\[\pm2\] is it right???

hero (hero):

I didn't see the square root..... wait..... Yes it is the sqrt (3x+1) f(g(1)) = sqrt (4) = 2

OpenStudy (anonymous):

\[ the answer is \pm 2 or the answer is +2... i dont know which one is \right??\]

OpenStudy (anonymous):

thanks:)..

OpenStudy (anonymous):

i got my answer is +/- 2 but my teacher said it's just +2... not -2

hero (hero):

Oh yeah, you can't take the sqrt of a negative number

hero (hero):

I retract my previous statement

hero (hero):

Sorry for the confusion, it's early morning here for me

OpenStudy (anonymous):

but (-2)^2= 4 too.

hero (hero):

That's true, but, hmmm. How can I explain this....

hero (hero):

There's a difference between squaring and taking a square root.

OpenStudy (anonymous):

so i dont know which one is the xactly answer +/- 2 or +2...???

hero (hero):

You can square two opposite numbers and get the same number, but you can only square root a positive number

hero (hero):

It's definitely +2 for the answer you seek.

hero (hero):

Once again, you cannot square root a negative number.....It's a mathematical law

OpenStudy (anonymous):

ok... thanks for helping :)

hero (hero):

You're welcome